mean height of 133 people is 155 cm stadard deviation is 15 how many persons are taller than 200cm


1
Expert's answer
2021-07-26T11:08:03-0400

Given,

"n=133\\Mean\ \ \mu=155\ cm\\SD\ \ \sigma=15\\X=200\ cm"


Using Normal Distribution:


"z=\dfrac{X-\mu}{\sigma/\sqrt n}=\dfrac{200-155}{15/\sqrt{133}}=34.5976"


So, number of people who are taller than 200 cm is equal to

"P(X>200)=P(z>34.5976)\approx 0"


hence, from the population of 133 people, No one is taller than 200 cm